1. A square wire loop 10cm on a side is oriented with its plane perpendicular to a magnetic field. The resistance of the loop is 5.0. How rapidly should the magnetic field change if a current of 2.0 A is to flow in the loop?
1
Expert's answer
2021-06-10T10:22:20-0400

According to Ohm's law, find the voltage:


"V=IR."

Using Faraday's law, express the voltage in terms of change of magnetic flux:


"V=\frac{\Delta \Phi}{\Delta t}=\frac{A\Delta B}{\Delta t},\\\space\\
\frac{\Delta B}{\Delta t}=\frac VA=\frac{IR}{a^2}=1000\text{ T/s}."